OpenStudy (anonymous):

A building lot in a city is shaped as a 30 -60 -90 degress triangle . the side oppisite the 30 angle messuses 41 feet a. Find the length of the side of the lot opposite the 60sa026-5.jpg angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 60degres angle in the lot. Write your answers as decimals rounded to four decimal places. (7 points)

OpenStudy (callisto):

 For part a, find (a) tan 30 = 41/(a) \[1/\sqrt3 = 41/ (a)\]\[(a) = 41\sqrt3\] For part (b), find (b) sin30 = 41 / (b) 1/2 = 41 /(b) b = 41x2 = 82 Now, you have all the sides and angle. Can you do part c yourself?